Operations Report – January 06th


Operations Report

Crew 219 Operations Report 06-JAN-2020

Sol: 1

Name of person filing report: Alejandro Luis Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 14

From what time last night: 17:30

To what time this morning: 07:30

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 49%

Diesel Reading – 73%

Station Propane Reading- 50%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 55

Water Meter (units)- 0147407,8

Water (static tank) (gal) – 513

Static to loft Pump used – Yes

Water in Green Hab (gal): 213.25

Water in Science Dome (gal): 0

Toilet Tank Emptied – Yes

Diemos rover used: Still in the Work shop


Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director


Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

EVA #1

Hours: 123.1

Beginning Charge: 100%

Ending Charge: 88%

EVA #2


Beginning Charge: 100%

Ending Charge: 92%

Currently Charging: Yes

Opportunity Rover used: Still in workshop


Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

EVA #1 and 2

Hours: 129.0

Beginning Charge: 100%

Ending Charge: 91%

EVA #2

Hours: 129.1

Beginning Charge: 100%

Ending Charge: 95%

Currently Charging: Yes

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): Yes, Outpost used ATV #1, and #3

Reason for use: Used for maintenance.

Oil Added? No

# hours the ATV’s were used today: 0.5

Notes on ATVs: Honda battery was not responsive, it needed to be jumped in order for it to turn on. This is the third occasion.

Habcar used and why, where? Used by the Outpost team to storage and Post Office.

CrewCar used and why, where? No

General Notes and Comments: Spare heat tape being kept in the RAM. The extension that is connected to the current heat tape has been relocated to the inner plugs in the lower deck. Also, Astronomer used the rest of the Orange marking tape for his experiment.

Summary of the internet: limited intermittent dysconnectivity

Summary of Suit and Radios: Suits #5 and Suits #1 do not have air flow. One headset that was not operating.

Summary of Hab: Followed instructions given by Director Rupert for flushing the toilet. Filled and emptied the toilet 2 times and added digester. Now operating on normal toilet procedures. See attached requested photos below regarding the black seal of the toilet. Also, the feedline to the loft tank is still leaking when the pump is turned on. Teflon tape would help improve the seal.

Summary of Science Dome operations: Nominal

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: See HSO report

Questions, concerns, and requests to Mission Support: Request for more Orange Marking Tape and Teflon tape

Copyright © The Mars Society. All rights reserved. | Main Site